OpenStudy (anonymous):

Find the polynomial function with roots 1, -2, and 5. work please

OpenStudy (campbell_st):

will it will be P(x) =k(x -1)(x+2)(x-5) k is a constant just expand and collect like terms

OpenStudy (saifoo.khan):

What's written above^

OpenStudy (saifoo.khan):

@campbell_st , is "k" necessary?

OpenStudy (anonymous):

There are infinitely many polynomials with those three roots. The "k" accounts for that.

OpenStudy (campbell_st):

yes as the polynomial P(x) = 2x^2 + 10x + 12 = 2(x^2 +5x +6) = 2(x+2)(x+3)

OpenStudy (saifoo.khan):

Ah, perfect.